OpenStudy (lgbasallote):

LGBAINTEGRAL: \[\huge \int (2x^3+5)^4 dx\] is this integrable??

OpenStudy (lgbasallote):

i dont think i want this im changing the question lol

OpenStudy (shayaan_mustafa):

Hi lgbasallote :) How are you?

OpenStudy (shayaan_mustafa):

Do you know which function is integrable or which is not? Do you know?

OpenStudy (lgbasallote):

there i think that's better

OpenStudy (lgbasallote):

i dont think that last one could be solveable by simple techniques

OpenStudy (anonymous):

Use binomial and expand it, if you can

OpenStudy (lgbasallote):

wont integration by parts work?

OpenStudy (shayaan_mustafa):

It is integrable dear.

OpenStudy (anonymous):

And this is totally Integrable.

mathslover (mathslover):

@vishweshshrimali5 may help u

OpenStudy (shayaan_mustafa):

Don't use by parts. Choose an easy way.

OpenStudy (shayaan_mustafa):

I agree with Ishaan94

OpenStudy (lgbasallote):

what easier way? expansion?

OpenStudy (shayaan_mustafa):

yes. Expand.

OpenStudy (lgbasallote):

hmmm...

OpenStudy (lgbasallote):

yeah i guess if it's just expansion then i can do the rest from there

OpenStudy (lgbasallote):

and seriously... \[\int \frac{x}{2+\tan x}dx\] that's integrable??? hahaha

OpenStudy (vishweshshrimali5):

yes this is integrable

OpenStudy (anonymous):

By Parts, perhaps

OpenStudy (anonymous):

If there would be 1 in the numerator then it can be integrated bu there is x.. Ha ha ha..

OpenStudy (lgbasallote):

lol wow

OpenStudy (lgbasallote):

i dont wanna attempt it o.O

OpenStudy (vishweshshrimali5):

My answer is for ur original question

OpenStudy (lgbasallote):

:yeah i know haha

OpenStudy (vishweshshrimali5):

@lgbasallote Very true

OpenStudy (shayaan_mustafa):

I think it is not integrable dear.

OpenStudy (shayaan_mustafa):

This function is not integrable.

OpenStudy (vishweshshrimali5):

for ur next question I don't have any idea for this but this is what wolfram says

OpenStudy (anonymous):

I just lost my work. I can only think of by parts. I haven't tried any calculator, so I don't know if its integral is defined.

OpenStudy (vishweshshrimali5):

but wait for Ishaan sir may be he would have a better way to solve

OpenStudy (shayaan_mustafa):

So this is not integrable. Right?

OpenStudy (experimentx):

perhaps not

OpenStudy (shayaan_mustafa):

Not perhaps. It is ofcourse not.

OpenStudy (vishweshshrimali5):

yes it is not integrable

OpenStudy (experimentx):

\[ \int \sin x /x dx\] is not integrable ... somehow this integration seems same type.